OpenStudy (anonymous):

Government data shows that 26% of the civilian force have at least four years of college and that 15% of the labor force works as laborers or operators of machines or vehicles. Can you conclude that because (0.26)(0.15) = 0.039 about 4% of the labor force are college educated laborers or operators? No, because the events are not independent No, because the events are not mutually exclusive Yes, by the multiplication rule Yes, by conditional probabilities

OpenStudy (anonymous):

need some help on this?

OpenStudy (anonymous):

Just a bit. I think it might be the first one, but I'm not so sure.

OpenStudy (anonymous):

\(P(A\cup B=P(A)+P(B)-P(A\cap B)\) this is for A "or" B \(P(A\cap B)=P(A)P(B)\) this is for A "and" B if they are independent events if events are mutually exclusive then \(P(A\cup B)=P(A)+P(B)\)

OpenStudy (anonymous):

So ... it's mutually exclusive?

OpenStudy (anonymous):

no, it's written in a confusing sort of way. they do that to make it more difficult. what you need to do is to define the events. Define event A as the person is college educated and define B as the person is a laborer or operator. then what is given is that \(P(A)P(B) = P(A\cap B)\) which is independence but it's not given that they are independent.

OpenStudy (anonymous):

So by the multiplicity rule is the closest answer

OpenStudy (anonymous):

no. it's the first choice.

OpenStudy (anonymous):

don't hate, just get a good tutor. or come here. you can master this stuff!

OpenStudy (anonymous):

Math in general just takes a while to click. Thank you though

OpenStudy (anonymous):

there are things like that in life... just be persistent and your effort will often be rewarded!!!